OpenStudy (anonymous):

]Assuming x > 0, which of these expressions is equivalent to 11 times the square root of 245 x to the third plus 9 times the square root of 45 x to the third?

OpenStudy (anonymous):

\[11\sqrt{245x ^{3}}+9\sqrt{45x ^{3}}\]

OpenStudy (anonymous):

\[5x \sqrt{104x} 20\sqrt{290x ^{6}} 20x \sqrt{290x} 104x \sqrt{5x}\]

OpenStudy (anonymous):